# Fermionic spinon theory of the hourglass spin excitation spectrum of the cuprates

### v1

@@ 1 Introduction
There have been a number of earlier studies suggesting the relevance of fractionalized spinon excitations (which carry spin 1/2 and electrical charge 0) to square lattice antiferromagnets [Hayden10, Ronnow15, Becca18, Ma2018, Yasir19, Ronnow25]. Recent work [Mascot2022, PNAS_pseudo1, Christos2024, BCS24, Sayantan25], reviewed in Ref. [Boulder25], has argued that such spinons are the key to understanding a number of observations in the underdoped cuprates. These arguments rely on the fractionalized Fermi liquid (FL*) theory [FLS1, Si2004, FLS2, APAV04, GroverSenthil10, Bonderson16, Tsvelik16, Vojta20, ElseSenthil, Tsvelik24] of the pseudogap phase: the FL* state of the doped square lattice antiferromagnet [Kaul07, ACL08, Qi10, Punk15, Zhang2020, Mascot2022] has 4 Fermi pockets of spin-1/2 charge $e$ quasiparticles, each of fractional area $p/8$, so that the total density of mobile quasiparticles is $p$. Such a state does not obey the Luttinger Fermi surface area [Luttinger60], but does satisfy the Oshikawa anomaly [MO00] by the presence of additional spinon excitations [FLS1, FLS2, APAV04, Bonderson16, ElseSenthil].
@@
Compelling evidence for such a FL* state comes from recent angle-dependent magnetoresistance (ADMR) measurements [fang_admr_2022, chan_yamaji_2024]. The ADMR displays significant evidence for the presence of hole pockets, and the Yamaji effect is consistent with the FL* pocket area of $p/8$ [Zhao_Yamaji_25]. Moreover, the Yamaji effect requires coherent inter-layer electronic transport, which is present for FL*, and unlikely for other pseudogap candidates [Zhao_Yamaji_25]. Other recent arguments for the FL* description of the pseudogap are summarized in Section 4.
@@
In this paper, we turn to the role of spinons in the spin-fluctuation spectrum, where they have a direct role. Our focus is on the well-studied ‘hourglass’ spectrum of the stripe state observed by neutron scattering in the cuprates near hole doping $p=1/8$ [Hayden2004, Vignolle2007, Chan2016].
@@
We use the $\pi$-flux spin liquid theory for the underlying square lattice antiferromagnet [AM88, Affleck-SU2, Fradkin88]. This has fermionic spinon excitations which carry spin $S=1/2$ and a dispersion with 2 massless Dirac nodes in the square lattice Brilloin zone. The spinons are coupled to an emergent SU(2) gauge field (the gauge SU(2) is distinct from, and commutes with, the spin rotation SU(2)). Fluctuations of the SU(2) gauge field in the insulator are expected to lead to confinement of the spinons into either the Néel state or the valence bond solid (VBS) state, but there is evidence for a deconfined critical state of the spinons between the Néel and VBS states [NRSS89prl, Wang17, Fuzzy24]. The free spinon dynamic structure factor is reviewed in Section 2.1, and this is in good correspondence with numerical studies of the Néel-VBS transition [Becca18]. The free spinon theory exhibits significant low energy spectral weight near momenta $(\pi,\pi)$, $(\pi,0)$, and $(0,0)$; Ref. [BCS24] argued that SU(2) gauge fluctuations strongly suppress the spectral weight near momenta $(0,0)$ and $(\pi,0)$.
@@
Resonant inelastic X-ray scattering (RIXS) experiments have investigated the high energy spin fluctuation spectrum in the cuprates [Keimer11, Hayden19, Keimer25]. Ref. [SSZaanen] has noted that the broad continuum scattering between momenta $(0,0)$ and $(\pi,0)$ bear a strong resemblance to the continuum spectral weight in the frustrated square lattice antiferromagnet near the Néel-VBS transition [Becca18]. The energy scales match well, and the main difference is that the insulating antiferromagnet has gapless spectral weight at $(\pi,0)$, while the doped cuprate has little spectral weight below $\sim 200$ meV. This difference could be a consequence of the doped carriers, whose influence will be studied here.
@@
The present paper will study a different route to confinement of the spinons at low energies, appropriate to the doped cuprates [PNAS_pseudo1]. We will assume a charge density wave (CDW) with static long-range order which is produced by the condensation of a Higgs field $B$ [BCS24]. This Higgs field carries a fundamental SU(2) gauge charge, and so its condensation gaps out all components of the SU(2) gauge field. The resulting $B$-condensed phase is ‘topologically trivial’, but nevertheless, signatures of the deconfined spinons can emerge at higher energies. And at low energies, as we shall demonstrate, such a spinon theory is able to reproduce the ‘hourglass’ spectrum of the CDW (‘stripe’) state. Our approach differs from other theories of the hourglass [Vojta2004, Vojta2006, Vojta2012, Vojta2013] by using spin $S=1/2$ excitations in a background of charge order, whereas the earlier works considered spin $S=1$ paramagnon excitations.
@@
We study the spin susceptibility of the underdoped cuprates using the Ancilla Layer Model (ALM) studied extensively in previous works [Zhang2020, Zhang2020_2, Nikolaenko2023, Christos2024, BCS24, Boulder25]. The ALM offers a unified description of both the FL* pseudogap metal with small Fermi surface, and a conventional Fermi liquid with large Fermi surface. It also hosts fractionalized spinon excitations that ensure consistency with the Oshikawa anomaly. In the following sections, we examine how these excitations affect the spin structure factor and make predictions for future neutron scattering experiments.
@@ 2 Theory
We start this section by formulating the ALM. The model has three layers: the first level consists of electrons described by tight-binding model, while the second and the third layers are filled with spins with Heisenberg interaction between them. The Hamiltonian of the ALM is [Zhang2020, Mascot2022, Christos2024]:
$H=\sum_{\bm{i},\bm{j},\alpha}\left\{-t_{\bm{i}\bm{j}}c^{\dagger}_{\bm{i},\alpha}c_{\bm{j},\alpha}+J_{K}{\bm{S}}_{c,\bm{i}}\cdot{\bm{S}}_{1,\bm{i}}+J_{1}{\bm{S}}_{1,\bm{i}}\cdot{\bm{S}}_{1,\bm{j}}+J_{\perp}{\bm{S}}_{1,\bm{i}}\cdot{\bm{S}}_{2,\bm{i}}+J_{2}{\bm{S}}_{2,\bm{i}}\cdot{\bm{S}}_{2,\bm{j}}\right\}\,$, (1)
where $\bm{i},\bm{j}\in\Lambda$, with $\Lambda=\{(x,y)\,|\,x,y\in\mathbb{Z}\}$ being a two-dimensional square lattice. For the first layer we employ the tight-binding parameters $t=0.22$ (nearest neighbor hopping), $t^{\prime}=-0.034$ (second neighbor hopping), $t^{\prime\prime}=0.036$ (third neighbor hopping), $t^{\prime\prime\prime}=-0.007$ (fourth neighbor hopping), and a chemical potential $\mu_{c}=-0.24$ (all in units of eV). These values have been extracted by fitting photoemission data from Ref. [He2011] in the overdoped regime, where large Fermi surface is clearly visible.
We use the Schwinger-fermion representation for the second ($a=1$) and third ($a=2$) layer spins: ${\bm{S}}_{a,\bm{i}}=f_{a,\bm{i},\alpha}^{\dagger}{\bm{\sigma}}_{\alpha\beta}f_{a,\bm{i},\beta}$ with the imposed constraint $\sum_{\alpha}f_{a,\bm{i},\alpha}^{\dagger}f_{a,\bm{i},\alpha}=1$, satisfied at every lattice site $i$. Throughout the paper, we adopt a mean-field approach and neglect all gauge field fluctuations emerging from the redundancy of the parton construction. Within the mean-field theory, the interacting terms in the Hamiltonian simplify to:
@@
where $\Phi\propto J_{K}\langle f_{1,\bm{i},\alpha}^{\dagger}c_{\bm{i},\alpha}\rangle\in\mathbb{R}$ and $B_{1,\bm{i}}\propto J_{\perp}\langle f_{1,\bm{i},\alpha}^{\dagger}f_{2,\bm{i},\alpha}\rangle$, $B_{2,\bm{i}}\propto J_{\perp}\langle f_{1,\bm{i},\alpha}^{\dagger}\varepsilon_{\alpha\beta}f^{\dagger}_{2,\bm{i},\beta}\rangle$, with $\varepsilon_{\alpha\beta}$ the unit antisymmetric tensor. The second layer is characterized by the hopping parameters $t_{1}=0.1,t_{1}^{\prime}=-0.03,t_{1}^{\prime\prime}=-0.01$, a chemical potential $\mu_{f}=0.009$, and a hybridization with the first layer $\Phi=0.09$. The finite hybridization between the first and second layers leads to the emergence of four Fermi pockets in the Brillouin zone with area $p/8$. Moreover, the intensity at back side of the pockets is subdued, which is consistent with Fermi arcs features observed in photoemission experiments. The parameters were chosen to reproduce the photoemission data for Bi2201 at a doping level of $p=0.206$ [Mascot2022, He2011]. The $B$ field would be discussed later, to incorporate the effects of CDW order. For now we assume $B_{a,\bm{i}}=0$.
A nonzero value of $\Phi$ reconstructs the top layer large Fermi surface, of (hole) volume $1+p$, into four hole pockets of total volume $p$. Because no translation symmetry breaking is occurring at this point, this is at odds with the conventional formulation of Luttinger’s theorem. To account for its violation, we assume the presence of a quantum spin liquid in the bottom layer, hosting fractionalized spinon excitations that account for the missing Luttinger volume [MO00, FLS1, FLS2]. In particular, we choose a $\pi$-flux spin liquid Ansatz of the form:
@@
with $e_{\bm{i},\bm{j}}=-e_{\bm{j},\bm{i}}$ and in the chosen gauge $e_{\bm{i},\bm{i}+\hat{\bm{x}}}=1$, $e_{\bm{i},\bm{i}+\hat{\bm{y}}}=(-1)^{x}$. We also set $t^{f_{2}}=0.14\,eV$ similar to Ref. [Christos2024]. The chosen gauge is not manifestly translation invariant in the $x$-direction. In fact, translations along the $x$-axis are implemented projectively $f_{2,\bm{i},\alpha}\to(-1)^{y}f_{2,\bm{i},\alpha}$, implying that all gauge invariant quantities computed with Hamiltonian (3) will be translation invariant. The Hamiltonian $H_{f_{2},f_{2}}$ can be diagonalized in the momentum space. We consider a supercell with two sites ($\delta_{\bm{i}}=0,1$) in the $x$-direction. After Fourier transform $f_{2,\delta_{i}}(\bm{r})=e^{i\bm{k}\bm{r}}f_{2,\delta_{i}}(\bm{k})$ the Hamiltonian in the basis $\{f^{\dagger}_{2,\delta_{i}}(\bm{k})\}$ is $2\times 2$ matrix with with spectrum $E_{\pm}(k)=\pm 2t^{f_{2}}\sqrt{\sin(k_{x})^{2}+\sin(k_{y})^{2}}$ and $\bm{k}$ is defined in the reduced Brillouin zone $\mathrm{BZ}^{\prime}=\{(k_{x},k_{y})\,|\,k_{x}\in(-\pi/2,\pi/2),\,k_{y}\in(-\pi,\pi)\}$. The spectrum features two Dirac cones (per spin projection $\alpha$) at $(k_{x},k_{y})=(0,0)$ and $(k_{x},k_{y})=(0,\pi)$.
@@ 2.2 Bare susceptibilities in the presence of charge density wave order
We now consider the susceptibilities when $B_{a,\bm{i}}\neq 0$. As found in Ref. [PNAS_pseudo1], the condensation of the field $B_{a,\bm{i}}$ has both effects of fully Higgsing the SU(2) gauge group associated with the Schwinger fermion decomposition in the bottom layer, and breaking some physical global symmetries, such as translations, time reversal, or U(1) charge conservation, depending on the spatial pattern of $B_{a,\bm{i}}$. In the following, we choose an Ansatz for $B_{a,\bm{i}}$ that only breaks translation symmetry along the $x$-axis:
@@
According to [PNAS_pseudo1], the CDW order parameter is given by:
$\rho_{i}=B_{i}^{\dagger}B_{i}=2b^{2}[1+(-1)^{x}\sin 2\theta][\cos\left(K_{x}x+\phi\right)+1]\,.$ (8)
One can check that the above Ansatz gives vanishing current and superconducting order parameters, $\Re[B^{\dagger}_{\bm{i}}e_{\bm{i},\bm{j}}B_{\bm{j}}]=0$, $B_{a,\bm{i}}\varepsilon_{ab}e_{\bm{i},\bm{j}}B_{b,\bm{j}}=0$. In the following, we restrict ourselves to a period-4 bond-centered CDW, given by $K_{x}=\frac{\pi}{2}$, $\phi=\frac{\pi}{4}$.
@@
$H_{int}=\sum_{ij}J_{ij}(1+gQ^{b}_{ij})f^{\dagger}_{i\alpha}f_{j\alpha}f^{\dagger}_{j\beta}f_{i\beta}=\sum_{ij}\tilde{J}_{ij}f^{\dagger}_{i\alpha}f_{j\alpha}f^{\dagger}_{j\beta}f_{i\beta},$ (13)
where $J_{ij}$ is a Heisenberg coupling. We will include RPA corrections from this interaction, following the application of such methods to the triangular lattice antiferromagnet [Batista18, LiZhang20, Knolle25]. After using the Fourier representation with $J_{ij}=\sum_{q}J_{\delta_{i}\delta_{j}}(\bm{q})e^{i\bm{q}(i-j)}$ and summing up the series, the RPA result in the spin-spin channel in matrix notation reads:

### v2

@@ Abstract
The Abdus Salam International Centre for Theoretical Physics, Strada Costiera 11, I-34151, Trieste, Italy.
@@ 1 Introduction
There have been a number of earlier studies suggesting the relevance of fractionalized spinon excitations (which carry spin 1/2 and electrical charge 0) to square lattice antiferromagnets [1, 2, 3, 4, 5, 6]. Recent work [7, 8, 9, 10, 11], reviewed in Refs. [12, 13], has argued that such spinons are the key to understanding a number of observations in the underdoped cuprates. These arguments rely on the fractionalized Fermi liquid (FL*) theory [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] of the pseudogap phase: the FL* state of the doped square lattice antiferromagnet [24, 25, 26, 27, 28, 7] has 4 Fermi pockets of spin-1/2 charge $e$ quasiparticles, each of fractional area $p/8$, so that the total density of mobile quasiparticles is $p$. Such a state does not obey the Luttinger Fermi surface area [29], but does satisfy the Oshikawa anomaly [30] by the presence of additional spinon excitations [14, 16, 17, 19, 22].
@@
Compelling evidence for such a FL* state comes from recent angle-dependent magnetoresistance (ADMR) measurements [31, 32]. The ADMR displays significant evidence for the presence of hole pockets which can tunnel coherently between square lattice layers, and the Yamaji effect is consistent with the FL* pocket area of $p/8$ [33]. Moreover, the Yamaji effect requires coherent inter-layer electronic transport, which is present for FL*, and unlikely for other pseudogap candidates [33]. Other recent arguments for the FL* description of the pseudogap are summarized in Section 4.
@@
In this paper, we turn to the role of spinons in the spin-fluctuation spectrum, where they have a direct role. Our focus is on the well-studied ‘hourglass’ spectrum of the stripe state observed by neutron scattering in the cuprates near hole doping $p=1/8$ [34, 35, 36]. The name ‘hourglass’ comes from the shape of the spectrum, which has a distinctive narrowing in the middle and widening at higher and lower energies.
@@
We use the $\pi$-flux spin liquid theory for the underlying square lattice antiferromagnet [37, 38, 39]. This has fermionic spinon excitations which carry spin $S=1/2$ and a dispersion with 2 massless Dirac nodes in the square lattice Brilloin zone. The spinons are coupled to an emergent SU(2) gauge field (the gauge SU(2) is distinct from, and commutes with, the spin rotation SU(2)). Fluctuations of the SU(2) gauge field in the insulator are expected to lead to confinement of the spinons into either the Néel state or the valence bond solid (VBS) state, but there is evidence for a deconfined critical state of the spinons between the Néel and VBS states [40, 41, 42]. The free spinon dynamic structure factor is reviewed in Section 2.1, and this is in good correspondence with numerical studies of the Néel-VBS transition [3]. (Ref. [3] employs a $\mathbb{Z}_{2}$ spin liquid for their numerics, but this spin liquid is closely connected to the SU(2) spin liquid [43], and has a similar spinon spectrum.) The free spinon theory exhibits significant low energy spectral weight near momenta $(\pi,\pi)$, $(\pi,0)$, and $(0,0)$; Ref. [10] argued that SU(2) gauge fluctuations strongly suppress the spectral weight near momenta $(0,0)$ and $(\pi,0)$.
@@
Resonant inelastic X-ray scattering (RIXS) experiments have investigated the high energy spin fluctuation spectrum in the cuprates [44, 45, 46]. Ref. [47] has noted that the broad continuum scattering between momenta $(0,0)$ and $(\pi,0)$ bear a strong resemblance to the continuum spectral weight in the frustrated square lattice antiferromagnet near the Néel-VBS transition [3]. The energy scales match well, and the main difference is that the insulating antiferromagnet has gapless spectral weight at $(\pi,0)$, while the doped cuprate has little spectral weight below $\sim 200$ meV. This difference could be a consequence of the doped carriers, whose influence will be studied here.
@@
The present paper will study a different route to confinement of the spinons at low energies, appropriate to the doped cuprates [8]. We will assume a charge density wave (CDW) with static long-range order which is produced by the condensation of a Higgs field $B$ [10]. This Higgs field carries a fundamental SU(2) gauge charge, and so its condensation gaps out all components of the SU(2) gauge field. The resulting $B$-condensed phase is ‘topologically trivial’, but nevertheless, signatures of the deconfined spinons can emerge at higher energies. And at low energies, as we shall demonstrate, such a spinon theory is able to reproduce the ‘hourglass’ spectrum of the CDW (‘stripe’) state. Our approach differs from other theories of the hourglass [48, 49, 50, 51] by using spin $S=1/2$ excitations in a background of charge order, whereas the earlier works considered spin $S=1$ paramagnon excitations.
@@
We study the spin susceptibility of the underdoped cuprates using the Ancilla Layer Model (ALM) studied extensively in previous works [28, 52, 53, 9, 10, 12]. The ALM offers a unified description of both the FL* pseudogap metal with small Fermi surface, and a conventional Fermi liquid with large Fermi surface. It also hosts fractionalized spinon excitations that ensure consistency with the Oshikawa anomaly. In the following sections, we examine how these excitations affect the spin structure factor and make predictions for future neutron scattering experiments.
@@ 2 Theory
We start this section by formulating the ALM. The model has three layers: the first level consists of electrons described by tight-binding model, while the second and the third layers are filled with spins with Heisenberg interaction between them. The Hamiltonian of the ALM is [28, 7, 9]:
$H=\sum_{\bm{i},\bm{j}}\left\{-t_{\bm{i}\bm{j}}c^{\dagger}_{\bm{i},\alpha}c_{\bm{j},\alpha}+J_{1,\bm{i}\bm{j}}\,{\bm{S}}_{1,\bm{i}}\cdot{\bm{S}}_{1,\bm{j}}+J_{2,\bm{i}\bm{j}}\,{\bm{S}}_{2,\bm{i}}\cdot{\bm{S}}_{2,\bm{j}}\right\}\,+\sum_{\bm{i}}\left\{J_{K}{\bm{S}}_{c,\bm{i}}\cdot{\bm{S}}_{1,\bm{i}}+J_{\perp}{\bm{S}}_{1,\bm{i}}\cdot{\bm{S}}_{2,\bm{i}}\right\}\,$, (1)
where $\bm{i},\bm{j}\in\Lambda$, with $\Lambda=\{(x,y)\,|\,x,y\in\mathbb{Z}\}$ being a two-dimensional square lattice. For the first layer we employ the tight-binding parameters $t=0.22$ (nearest neighbor hopping), $t^{\prime}=-0.034$ (second neighbor hopping), $t^{\prime\prime}=0.036$ (third neighbor hopping), $t^{\prime\prime\prime}=-0.007$ (fourth neighbor hopping), and a chemical potential $\mu_{c}=-0.24$ (all in units of eV). As in Ref. [7], these values have been extracted by fitting photoemission data of Ref. [54] in the overdoped regime, where a large Fermi surface is clearly visible. Indeed, in the overdoped phase, the second and third layers of the ALM effectively decouple from the rest of the system by forming rung singlets. Thus, the physics is described by the first layer only, which is a Fermi liquid with a large Fermi surface, whose shape is determined by the hoppings $t_{\bm{i}\bm{j}}$, as in conventional tight-binding models.
We use the Schwinger-fermion representation for the second ($a=1$) and third ($a=2$) layer spins: ${\bm{S}}_{a,\bm{i}}=(1/2)f_{a,\bm{i},\alpha}^{\dagger}{\bm{\sigma}}_{\alpha\beta}f_{a,\bm{i},\beta}$ with the imposed constraint $\sum_{\alpha}f_{a,\bm{i},\alpha}^{\dagger}f_{a,\bm{i},\alpha}=1$, satisfied at every lattice site $i$. Throughout the paper, we adopt a mean-field approach and neglect all gauge field fluctuations emerging from the redundancy of the parton construction. Within the mean-field theory, the interacting terms in the Hamiltonian simplify to:
@@
where $\Phi\propto J_{K}\langle f_{1,\bm{i},\alpha}^{\dagger}c_{\bm{i},\alpha}\rangle\in\mathbb{R}$ and $B_{1,\bm{i}}\propto J_{\perp}\langle f_{1,\bm{i},\alpha}^{\dagger}f_{2,\bm{i},\alpha}\rangle$, $B_{2,\bm{i}}\propto J_{\perp}\langle f_{1,\bm{i},\alpha}^{\dagger}\varepsilon_{\alpha\beta}f^{\dagger}_{2,\bm{i},\beta}\rangle$, with $\varepsilon_{\alpha\beta}$ the unit antisymmetric tensor. The second layer is characterized by the hopping parameters $t_{1}=0.1,t_{1}^{\prime}=-0.03,t_{1}^{\prime\prime}=-0.01$, a chemical potential $\mu_{f}=0.009$, and a hybridization with the first layer $\Phi=0.09$. The finite hybridization between the first and second layers leads to the emergence of four Fermi pockets in the Brillouin zone with area $p/8$. Moreover, the intensity at back side of the pockets is subdued, which is consistent with Fermi arcs features observed in photoemission experiments. The parameters were chosen to reproduce the photoemission data for Bi2201 at a doping level of $p=0.206$ [7, 54]. The $B$ field would be discussed later, to incorporate the effects of CDW order. For now we assume $B_{a,\bm{i}}=0$.
A nonzero value of $\Phi$ reconstructs the top layer large Fermi surface, of (hole) volume $1+p$, into four hole pockets of total volume $p$. Because no translation symmetry breaking is occurring at this point, this is at odds with the conventional formulation of Luttinger’s theorem. To account for its violation, we assume the presence of a quantum spin liquid in the bottom layer, hosting fractionalized spinon excitations that account for the missing Luttinger volume [30, 14, 16]. In particular, we choose a $\pi$-flux spin liquid Ansatz of the form:
@@
with $e_{\bm{i},\bm{j}}=-e_{\bm{j},\bm{i}}$ and in the chosen gauge $e_{\bm{i},\bm{i}+\hat{\bm{x}}}=1$, $e_{\bm{i},\bm{i}+\hat{\bm{y}}}=(-1)^{x}$. We also set $t^{f_{2}}=0.14\,eV$ similar to Ref. [9]. The chosen gauge is not manifestly translation invariant in the $x$-direction. In fact, translations along the $x$-axis are implemented projectively $f_{2,\bm{i},\alpha}\to(-1)^{y}f_{2,\bm{i},\alpha}$, implying that all gauge invariant quantities computed with Hamiltonian (3) will be translation invariant. The Hamiltonian $H_{f_{2},f_{2}}$ can be diagonalized in the momentum space. We consider a supercell with two sites ($\delta_{\bm{i}}=0,1$) in the $x$-direction. After Fourier transform $f_{2,\delta_{i}}(\bm{r})=e^{i\bm{k}\bm{r}}f_{2,\delta_{i}}(\bm{k})$ the Hamiltonian in the basis $\{f^{\dagger}_{2,\delta_{i}}(\bm{k})\}$ is $2\times 2$ matrix with with spectrum $E_{\pm}(k)=\pm 2t^{f_{2}}\sqrt{\sin(k_{x})^{2}+\sin(k_{y})^{2}}$ and $\bm{k}$ is defined in the reduced Brillouin zone $\mathrm{BZ}^{\prime}=\{(k_{x},k_{y})\,|\,k_{x}\in(-\pi/2,\pi/2),\,k_{y}\in(-\pi,\pi)\}$. The spectrum features two Dirac cones (per spin projection $\alpha$) at $(k_{x},k_{y})=(0,0)$ and $(k_{x},k_{y})=(0,\pi)$.
@@ 2.2 Bare susceptibilities in the presence of charge density wave order
We now consider the susceptibilities when $B_{a,\bm{i}}\neq 0$. As found in Ref. [8], the condensation of the field $B_{a,\bm{i}}$ has both effects of fully Higgsing the SU(2) gauge group associated with the Schwinger fermion decomposition in the bottom layer, and breaking some physical global symmetries, such as translations, time reversal, or U(1) charge conservation, depending on the spatial pattern of $B_{a,\bm{i}}$. In the following, we choose an Ansatz for $B_{a,\bm{i}}$ that only breaks translation symmetry along the $x$-axis:
@@
According to [8], the CDW order parameter is given by:
$\rho_{\bm{i}}=B_{\bm{i}}^{\dagger}B_{\bm{i}}=2b^{2}[1+(-1)^{x}\sin 2\theta][\cos\left(K_{x}x+\phi\right)+1]\,.$ (8)
One can check that the above Ansatz gives vanishing current and superconducting order parameters, $\Re[B^{\dagger}_{\bm{i}}e_{\bm{i},\bm{j}}B_{\bm{j}}]=0$, $B_{a,\bm{i}}\varepsilon_{ab}e_{\bm{i},\bm{j}}B_{b,\bm{j}}=0$. In the following, we restrict ourselves to a period-4 bond-centered CDW, given by $K_{x}=\frac{\pi}{2}$, $\phi=\frac{\pi}{4}$. Such CDW order is well established in some cuprate materials, for example optimally doped LBCO [55, 56].
@@
$H_{int}=\sum_{\bm{i}\bm{j}}J_{2,\bm{i}\bm{j}}(1+gQ^{b}_{\bm{i}\bm{j}})f^{\dagger}_{\bm{i}\alpha}f_{\bm{j}\alpha}f^{\dagger}_{\bm{j}\beta}f_{\bm{i}\beta}=\sum_{\bm{i}\bm{j}}\tilde{J}_{\bm{i}\bm{j}}f^{\dagger}_{\bm{i}\alpha}f_{\bm{j}\alpha}f^{\dagger}_{\bm{j}\beta}f_{\bm{i}\beta},$ (13)
where $J_{2,\bm{i}\bm{j}}$ are the Heisenberg couplings in the bottom layer. We choose $J_{2,\bm{i}\bm{j}}$ so that its Fourier transform $J_{2}(\bm{q})$ is given by
$\begin{split}J_{2}(\bm{q})=&2J\left(\cos q_{x}+\cos q_{y}\right)+4J^{\prime}\cos q_{x}\cos q_{y}+2J^{\prime\prime}\left(\cos 2q_{x}+\cos 2q_{y}\right)\\ &+4J^{\prime\prime\prime}\left(\cos 2q_{x}\,\cos q_{y}+\cos q_{x}\,\cos 2q_{y}\right)\,.\end{split}$ (14)
We will include RPA corrections from this interaction, following the application of such methods to the triangular lattice antiferromagnet [57, 58, 59]. After using the Fourier representation
$J_{ij}=\sum_{q}J_{\delta_{i}\delta_{j}}(\bm{q})e^{i\bm{q}(i-j)}\,$, (15)
and summing up the series, the RPA result in the spin-spin channel in matrix notation reads: